OpenStudy (anonymous):

Solve t = p+3r over 6 for r

OpenStudy (lgbasallote):

\[t = p + \frac{3r} 6\] is that the question??

OpenStudy (anonymous):

no p+3r are together

OpenStudy (lgbasallote):

\[t = \frac{p+3r} 6\] okay...first multiply both sides by 6..what do you get?

OpenStudy (anonymous):

t x 6 = p+3r ?

OpenStudy (lgbasallote):

right. no subtract p from both sides

OpenStudy (anonymous):

i'm kind of confused is it t x 6 - p =3r ?

OpenStudy (lgbasallote):

yep..but rewrite t*6 as 6t...so it's neater

OpenStudy (lgbasallote):

6t - p = 3r now divide both sides by 3

OpenStudy (anonymous):

okay and is it 6t-p/3 = R as the answer

OpenStudy (lgbasallote):

right

OpenStudy (anonymous):

Thanks

OpenStudy (lgbasallote):

welcome